An ideal gas, initially at P1 = 0.75 bar, undergoes an adiabatic expansion from V1 = 10.0 L to V2 = 30.0 L. What is the final pressure and temperature of this gas? Explain how you got this.
Answer #1

Here we are assuming that the ideal gas is monoatomic,

gamma = 5/3

For adiabatic expansion,

PV^gamma = constant

0.75*10^(5/3) = P2 *30^(5/3)

P2 = 0.12 bar answer

Now assuming that number of moles taken was 1 mol (this will be needed as initial temperature was not given). Now using ideal gas equation,

T2 = P2V2/nR = 0.12*10^5*30*10^-3/[1*8.3]

= 43.37 K